A. Given the following information about the hyperbola, find both its standard and general form:

1. FOCI: (-4, 0) (4, 0); VERTICES: (-3, 0) (3, 0)

2. FOCI: (-3, -3) (-3, 13); VERTICES: (-3, 0) (-3, 10)


B. Given the standard form of the equation of hyperbola, find the following, then sketch the graph:


A. orientation

B. opening of branches

C. center

D. transverse axis

E. conjugate axis

F. distance of the foci

G. vertices

H. foci

I. co-vertices

J. asymptotes


1.(x+4)^2/9 − (y+3)^2/4 = 1


2.(y+2)^2/36 − (x−1)^2/49 = 1


C. Convert each general form to standard form of the equation of hyperbola: (10 points)

1. 9y^2 − 4x^2 − 18y + 24x − 63 = 0


2. 9x^2 - 4y^2 - 90x + 32y - 163 = 0





1
Expert's answer
2021-03-18T10:06:17-0400

A. 1. Basic “horizontal” hyperbola:

Equation: "\dfrac{x^2}{a^2 }-\dfrac{y^2 }{b^2 }=1"

"c^2=a^2+b^2"

Foci: "(-c, 0), (c, 0)"

Vertices: "(-a, 0), (a, 0)"

"a=3, c=4"


"b^2=4^2 -3^2=7"

The standard form of an equation of a hyperbola


"\dfrac{x^2}{9 }-\dfrac{y^2 }{7 }=1"

The general form of an equation of a hyperbola


"6x^2-9y^2-63=0"

2. The center is "(h, k)" and the vertices are "(h, k\pm b)"

Equation: "\dfrac{(y-k)^2}{b^2 }-\dfrac{(x-h)^2 }{a^2 }=1"

Foci: "(h, k-c), (h, k+c)"

"h=-3, b=5, c=8, k=5"


"a^2=8^2-5^2=39"

The standard form of an equation of a hyperbola


"\dfrac{(y-5)^2}{25}-\dfrac{(x+3)^2}{39 }=1"

The general form of an equation of a hyperbola


"-25x^2-150x+39y^2-390y-225=0"

B.

1.


"\dfrac{(x+4)^2}{9}-\dfrac{(y+3)^2}{4 }=1"

A. Horizontal oriented hyperbola 

B. Opens left and right.

C. Center "(-4,-3)."

D. Major (transverse) axis length: 6

E. Minor (conjugate) axis length: 4

F. "c=\sqrt{13}"

Focal Parameter: "\dfrac{4\sqrt{13}}{13}"

G. Vertices: "(-7, -3), (-1, -3)"

H. Foci: "(-4-\sqrt{13}, -3), (-4+\sqrt{13}, -3)"

I. Co-vertices: "(-4, -5), (-4, -1)"

J. Asymptotes:"y=-\dfrac{2}{3}x-\dfrac{17}{3}, y=\dfrac{2}{3}x-\dfrac{1}{3}"


2.


"\dfrac{(y+2)^2}{36}-\dfrac{(x-1)^2}{49 }=1"

A. Verticaloriented hyperbola 

B. Opens upward and downward.

C. Center "(1,-2)."

D. Major (transverse) axis length: 12

E. Minor (conjugate) axis length: 14

F. "c=\sqrt{85}"

Focal Parameter: "\dfrac{49\sqrt{85}}{85}"

G. Vertices: "(1, -8), (1, 4)"

H. Foci: "(1,-2-\sqrt{85}), (1,-2+\sqrt{85})"

I. Co-vertices: "(-6, -2), (8, -2)"

J. Asymptotes:"y=-\dfrac{6}{7}x-\dfrac{8}{7}, y=\dfrac{6}{7}x-\dfrac{20}{7}"


C.

1.


"9y^2-4x^2-18y+24x-63=0"

"9(y^2-2y+1)-4(x^2-6x+9)=36"

"\dfrac{(y-1)^2}{4}-\dfrac{(x-3)^2}{9}=1"

2.


"9x^2-4y^2-90x+32y-163=0"

"9(x^2-10x+25)-4(y^2-8y+16)=324"

"\dfrac{(x-5)^2}{36}-\dfrac{(y-4)^2}{81}=1"